# Haoran Guan

### v1

@@ 4.2 Theoretical results of SLHC2 and general settings
4.4.1 Orthogonality
In the beginning, we proof the orthogonality of SSLHC3. With (85), we can have
$\Omega_{2}\Omega_{1}PXR_{ss}^{-1}=\Omega_{2}\Omega_{1}LUR_{ss}^{-1}+E_{LU}R_{ss}^{-1}.$ (143)
Based on (86)-(88), we can have
$\Omega_{2}\Omega_{1}PXR_{ss}^{-1}$ $=$ $\Omega_{2}\Omega_{1}LUR_{ss}^{-1}+E_{LU}R_{ss}^{-1}$ (144)
$=$ $(L_{ss}-E_{L_{ss}})UR_{ss}^{-1}+E_{LU}R_{ss}^{-1}$
$=$ $(W_{ss}G_{ss}+E_{HQ_{ss}}-E_{L_{ss}})UR_{ss}^{-1}+E_{LU}R_{ss}^{-1}$
$=$ $W_{ss}G_{ss}UR_{ss}^{-1}+E_{HQss}UR_{ss}^{-1}-E_{Lss}UR_{ss}^{-1}+E_{LU}R_{ss}^{-1}$
$=$ $W_{ss}(R_{ss}+E_{Rss})R_{ss}^{-1}+E_{HQss}UR_{ss}^{-1}-E_{Lss}UR_{ss}^{-1}+E_{LU}R_{ss}^{-1}$
$=$ $W_{ss}+W_{ss}E_{Rss}R_{ss}^{-1}+E_{HQss}UR_{ss}^{-1}-E_{Lss}UR_{ss}^{-1}+E_{LU}R_{ss}^{-1}.$
Based on (144), when $W_{ss}$ is orthogonal, we can have
$\|\Omega_{2}\Omega_{1}PXR_{ss}^{-1}-W_{ss}\|_{2}$$\leq$ $\|E_{Rss}\|_{2}\|R_{ss}^{-1}\|_{2}+\|E_{HQss}\|_{2}\|UR_{ss}^{-1}\|_{2}$ (145)
$+$ $\|E_{Rss}\|_{2}\|UR_{ss}^{-1}\|_{2}+\|E_{LU}\|_{2}\|R_{ss}^{-1}\|_{2}.$
Therefore, we put (56), (118), (124), (126) and (136) into (145) and based on (103) and (104), we can have
$\|\Omega_{2}\Omega_{1}PXR_{ss}^{-1}-W_{ss}\|_{2}$$\leq$ $1.13n^{2}\sqrt{n}{\bf u}\cdot\sqrt{1+\epsilon_{b}}\cdot(1+1.08s_{2}n\sqrt{n}{\bf u})\cdot\|L\|_{2}\|U\|_{2}$ (146)
$\cdot$ $\frac{1.12}{\sqrt{1-\epsilon_{s}}}\cdot\frac{1}{\sigma_{min}(L)\sigma_{min}(U)}+1.1s_{2}n^{2}{\bf u}\cdot\sqrt{1+\epsilon_{b}}\cdot\|L\|_{2}$
$\cdot$ $\frac{1.12}{\sigma_{min}(L)}\cdot\sqrt{1-\epsilon_{s}}+1.1\cdot\sqrt{1+\epsilon_{b}}\cdot s_{1}{\bf u}\cdot\sqrt{n}\|L\|_{2}$
$\cdot$ $\frac{1.12}{\sigma_{min}(L)}\cdot\sqrt{1-\epsilon_{s}}+1.1n^{2}{\bf u}\|L\|_{2}\|U\|_{2}$
@@ 4.4.1 Orthogonality
with probability at least $1-p_{f}$. Since $P$ is a permutation matrix, we can easily have
$\sigma_{i}(\Omega_{2}\Omega_{1}XR_{ss}^{-1})=\sigma_{i}(\Omega_{2}\Omega_{1}PXR_{ss}^{-1})$ (147)
when $i=1,2,\cdots,n$. Therefore, with (146) and (147), we can have
$0.8\leq\sigma_{min}(\Omega_{2}\Omega_{1}XR_{ss}^{-1})\leq\|\Omega_{2}\Omega_{1}XR_{ss}^{-1}\|_{2}\leq 1.2$ (148)
with probability at least $1-p_{f}$. Based on (14) and (15), we can have
$\frac{4}{5\sqrt{1+\epsilon_{b}}}\leq\sigma_{min}(XR_{ss}^{-1})\leq\|XR_{ss}^{-1}\|_{2}\leq\frac{6}{5\sqrt{1-\epsilon_{s}}}$ (149)
with probability at least $1-p_{f}$. With (89), we can have
$Q_{ss}-XR_{ss}^{-1}=E_{QRss}R_{ss}^{-1}.$ (150)
Based on (126) and (139), when (104) is satisfied, we can have
$\|Q_{ss}-XR_{ss}^{-1}\|_{2}$$\leq$ $\|E_{QRss}\|_{2}\|R_{ss}^{-1}\|_{2}$ (151)
$\leq$ $1.15n^{2}\sqrt{n}{\bf u}\cdot\sqrt{1+\epsilon_{b}}\cdot(1+1.08s_{2}n\sqrt{n}{\bf u})\cdot\|L\|_{2}\|U\|_{2}\|Q_{ss}\|_{2}$
@@
$\leq$ $0.06\|Q_{ss}\|_{2}$
with probability at least $1-p_{f}$. Based on Lemma 1 and (89), we can have
$\sigma_{min}(XR_{ss}^{-1})-\|Q_{ss}-XR_{ss}^{-1}\|_{2}\leq\sigma_{min}(Q_{ss})\leq\|Q_{ss}\|_{2}\leq\|XR_{ss}^{-1}\|_{2}+\|Q_{ss}-XR_{ss}^{-1}\|_{2}.$ (152)
Based on (149) and (151), we can get $\|Q_{ss}\|_{2}\leq\frac{6}{5\sqrt{1-\epsilon_{s}}}+0.06\|Q_{ss}\|_{2}$ with probability at least $1-p_{f}$. Therefore, it is easy to have
$\|Q_{ss}\|_{2}\leq\frac{1.28}{\sqrt{1-\epsilon_{s}}}$ (153)
with probability at least $1-p_{f}$. With (153), we can have $\|Q_{ss}-XR_{ss}^{-1}\|_{2}\leq\frac{0.08}{\sqrt{1-\epsilon_{s}}}$ with probability at least $1-p_{f}$. Therefore, with (149), we can have
$\sigma_{min}(Q_{ss})\geq\frac{4}{5\sqrt{1+\epsilon_{b}}}-\frac{0.08}{\sqrt{1-\epsilon_{s}}}$ (154)
with probability at least $1-p_{f}$. With (153) and (154), we can have
$\kappa_{2}(Q_{ss})$$\leq$ $\frac{\|Q_{ss}\|_{2}}{\sigma_{min}(Q_{ss})}$ (155)
@@
with probability at least $1-p_{f}$. Under (98), we can easily get $\frac{1.28}{0.8\cdot\sqrt{\frac{1-\epsilon_{s}}{1+\epsilon_{b}}}-0.08}\leq\frac{1}{8\sqrt{mn{\bf u}+n(n+1){\bf u}}}$. Then based on the theoretical results of error , we can get
@@
with probability at least $1-p_{f}$. Therefore, (99) is proved.
4.4.2 Residual
In the following, we focus on the residual. Regarding (92) and (93), we can have
$Q_{1ss}R_{2ss}$ $=$ $(Q_{ss}+E_{3ss})R_{1ss}^{-1}\cdot(R_{1ss}R_{ss}+E_{4ss})$ (156)
@@ 4.4.2 Residual
Therefore, with (156), we can have
$\|Q_{1ss}R_{2ss}-X\|_{F}\leq\|Q_{ss}R_{ss}-X\|_{F}+\|E_{3ss}\|_{F}\|R_{ss}\|_{2}+\|E_{4ss}\|_{F}\|Q_{1ss}\|_{2}.$ (157)
With Lemma 2, (92) and (93), we can have
$\|Q_{1ss}R_{2ss}-X\|_{F}$$\leq$ $\|Q_{ss}R_{ss}-X\|_{F}+\|E_{3ss}\|_{F}\|R_{ss}\|_{2}+\|E_{4ss}\|_{F}\|Q_{1ss}\|_{2}$ (158)
$\leq$ $1.1n^{2}{\bf u}(\|Q_{ss}\|_{2}\|R_{ss}\|_{2}+2\|Q_{1ss}\|_{2}\|R_{1ss}\|_{2}\|R_{ss}\|_{2}).$
For $\|Q_{1ss}R_{2ss}-X\|_{F}$, we can do it in a similar way with that of LUChol . Regarding $\|Q_{1ss}\|_{2}$, with theoretical results in error and (155), we can have
$\|Q_{1ss}^{\top}Q_{1ss}-I\|_{F}$$\leq$ $\frac{5}{64}\cdot(8\kappa_{2}(Q_{ss})\sqrt{mn{\bf u}+n(n+1){\bf u}})^{2}$ (159)
@@
with probability at least $1-p_{f}$. Therefore, we can get
$\|Q_{1ss}\|_{2}\leq\sqrt{1+h_{1}}$ (160)
with probability at least $1-p_{f}$. With corresponding results in error and (153), we can have
$\|R_{1ss}\|_{2}$$\leq$ $\frac{65}{63}\|Q_{ss}\|_{2}$ (161)
@@
with probability at least $1-p_{f}$. Based on (89) and (140), we can have
$\|Q_{ss}R_{ss}-X\|_{2}$$\geq$ $\|Q_{ss}R_{ss}\|_{2}-\|E_{QRss}\|_{2}$ (162)
$\geq$ $\sigma_{min}(Q_{ss})\|R_{ss}\|_{2}-1.1n^{2}\|Q_{ss}\|_{2}\|R_{ss}\|_{2}.$
@@
$\|R_{ss}\|_{2}\leq\frac{\|X\|_{2}}{\sigma_{min}(Q_{ss})-1.1n^{2}{\bf u}\|Q_{ss}\|_{2}}.$ (163)
We put (153) and (154) into (163) and based on (102), we can get
$\|R_{ss}\|_{2}$$\leq$ $\frac{\|X\|_{2}}{\frac{4}{5\sqrt{1+\epsilon_{b}}}-\frac{0.08}{\sqrt{1-\epsilon_{s}}}-1.1n^{2}{\bf u}\cdot\frac{1.28}{\sqrt{1-\epsilon_{s}}}}$ (164)
@@
with probability at least $1-p_{f}$. Here, (98) guarantees $\frac{4}{5\sqrt{1+\epsilon_{b}}}-\frac{0.11}{\sqrt{1-\epsilon_{s}}}>0$. Therefore, We put (153), (160), (161) and (164) into (157) and we can have
$\|Q_{1ss}R_{2ss}-X\|_{F}$$\leq$ $\|Q_{ss}R_{ss}-X\|_{F}+\|E_{3ss}\|_{F}\|R_{ss}\|_{2}+\|E_{4ss}\|_{F}\|Q_{1ss}\|_{2}$ (165)
@@
with probability at least $1-p_{f}$.
Regarding $\|Q_{2}^{\top}R_{4}-X\|_{F}$, with (96) and (97), we can have
$Q_{2ss}R_{4ss}$$\leq$ $(Q_{1ss}+E_{7ss})R_{3ss}^{-1}(R_{3ss}R_{2ss}+E_{8ss})$ (166)
@@
$\|Q_{2ss}R_{4ss}-X\|_{F}\leq\|Q_{1ss}R_{2ss}-X\|_{F}+\|E_{7ss}\|_{F}\|R_{4ss}\|_{2}+\|Q_{2ss}\|_{F}\|E_{8ss}\|_{F}.$ (167)
According to the existing results in error , we can have
$\|Q_{2ss}\|_{2}\leq 1.1$ (168)
with probability at least $1-p_{f}$. Moreover, based on (97), we can have
$\|R_{4ss}\|_{2}\leq\|R_{3ss}\|_{2}\|R_{2ss}\|_{2}+\|E_{8ss}\|_{2}.$ (169)
Similar to (169), regarding $\|R_{2ss}\|_{2}$, with (93), we can have
$\|R_{2ss}\|_{2}\leq\|R_{1ss}\|_{2}\|R_{ss}\|_{2}+\|E_{4ss}\|_{2}.$ (170)
With Lemma 2, (102), (161) and (164), we can estimate $\|E_{4ss}\|_{F}$ as
$\|E_{4ss}\|_{F}$$\leq$ $1.1n^{2}{\bf u}\|R_{1ss}\|_{2}\|R_{ss}\|_{2}$ (171)
@@
with probability at least $1-p_{f}$. Therefore, we put (161), (164) and (171) into (170) and we can get
$\|R_{2ss}\|_{2}$$\leq$ $\|R_{1ss}\|_{2}\|R_{ss}\|_{2}+\|E_{4ss}\|_{2}$ (172)
@@
with probability at least $1-p_{f}$. Similar to (161) and (160), we can bound $\|R_{3ss}\|_{2}$ as
$\|R_{3ss}\|$$\leq$ $\frac{65}{63}\|Q_{1ss}\|_{2}$ (173)
@@
with probability at least $1-p_{f}$. Therefore, with Lemma 2, (102), (172) and (173), we can estimate $\|E_{8ss}\|_{F}$ as
$\|E_{8ss}\|_{F}$$\leq$ $1.1n^{2}{\bf u}\|R_{3ss}\|_{2}\|R_{2ss}\|_{2}$ (174)
@@
with probability at least $1-p_{f}$. With Lemma 2, (168) and (173), we can bound $\|E_{7ss}\|_{F}$ as
$\|E_{7ss}\|_{F}$$\leq$ $1.1n^{2}{\bf u}\|Q_{2ss}\|_{2}\|R_{3ss}\|_{2}$ (175)
@@
with probability at least $1-p_{f}$. Moreover, with (102), we put (172)-(174) into (169) and we can estimate $\|R_{4ss}\|_{2}$ as
$\|R_{4ss}\|_{2}$$\leq$ $\|R_{3ss}\|_{2}\|R_{2ss}\|_{2}+\|E_{8ss}\|_{2}$ (176)
@@
with probability at least $1-p_{f}$. In all, we put (165), (168), (174)-(176) into (167) and we can get
@@
$\leq$ $[1.79(1+h_{1})+4.63\sqrt{1+h_{1}}+1.41]\cdot\frac{h_{2}}{\sqrt{1-\epsilon_{s}}}\cdot n^{2}{\bf u}\|X\|_{2}$
with probability at least $1-p_{f}$. (100) is proved. In all, Theorem 4.1 is proved.
4.5 Proof of Theorem 4.2
Proof
In the proof of Theorem 4.2, we focus directly on (159) and (165). When we consider the single-sketching case with $\Omega=\Omega_{2}$, we can set $\Omega_{1}\in\mathbb{R}^{m\times m}$ to be the identity matrix $I_{m}$, resulting in $s_{1}=m$. Consequently, if $\Omega_{1}\in\mathbb{R}^{s_{1}\times m}$ is a $(\epsilon_{1},p_{1},n)$ oblivious $l_{2}$-subspace embedding in $\mathbb{R}^{m}$, we can take $\epsilon_{1}=0$ and $p_{1}=0$. Next, when $\Omega\in\mathbb{R}^{s\times m}$ is a $(\epsilon,p,n)$ oblivious $l_{2}$-subspace embedding in $\mathbb{R}^{s_{1}}$, we let $\epsilon=\epsilon_{b}=\epsilon_{s}$, $p=p_{f}$ and $s=s_{2}$ in (159) and (164). With these settings, we can derive (106) and (107). Thus, Theorem 4.2 is proved.
5 Numerical experiments

### v2

@@ 4.2. Theoretical results of SLHC2 and general settings
4.4.1. Orthogonality
In the beginning, we proof the orthogonality of SSLHC3. With (4.1)-(4.4), we have
(4.56) $\Omega_{2}\Omega_{1}PXR_{ss}^{-1}$ $=$ $\Omega_{2}\Omega_{1}LUR_{ss}^{-1}+\Omega_{2}\Omega_{1}E_{LU}R_{ss}^{-1}$
$=$ $(L_{ss}-E_{L_{ss}})UR_{ss}^{-1}+\Omega_{2}\Omega_{1}E_{LU}R_{ss}^{-1}$
$=$ $(W_{ss}G_{ss}+E_{HQ_{ss}}-E_{L_{ss}})UR_{ss}^{-1}+\Omega_{2}\Omega_{1}E_{LU}R_{ss}^{-1}$
$=$ $W_{ss}G_{ss}UR_{ss}^{-1}+E_{HQss}UR_{ss}^{-1}-E_{Lss}UR_{ss}^{-1}+\Omega_{2}\Omega_{1}E_{LU}R_{ss}^{-1}$
$=$ $W_{ss}(R_{ss}+E_{Rss})R_{ss}^{-1}+E_{HQss}UR_{ss}^{-1}-E_{Lss}UR_{ss}^{-1}+\Omega_{2}\Omega_{1}E_{LU}R_{ss}^{-1}$
$=$ $W_{ss}+W_{ss}E_{Rss}R_{ss}^{-1}+E_{HQss}UR_{ss}^{-1}-E_{Lss}UR_{ss}^{-1}+\Omega_{2}\Omega_{1}E_{LU}R_{ss}^{-1}.$
Based on (4.56), when $W_{ss}$ is orthogonal, we can have
(4.57) $\|\Omega_{2}\Omega_{1}PXR_{ss}^{-1}-W_{ss}\|_{2}$$\leq$ $\|E_{Rss}\|_{2}\|R_{ss}^{-1}\|_{2}+\|E_{HQss}\|_{2}\|UR_{ss}^{-1}\|_{2}$
$+$ $\|E_{Lss}\|_{2}\|UR_{ss}^{-1}\|_{2}+\|\Omega_{1}\|_{2}\|\Omega_{2}\|_{2}\|E_{LU}\|_{2}\|R_{ss}^{-1}\|_{2}.$
Therefore, we put (3.40), (4.24), (4.31), (4.37), (4.39) and (4.49) into (4.57) and based on (4.13) and (4.14), we can have
(4.58) $\|\Omega_{2}\Omega_{1}PXR_{ss}^{-1}-W_{ss}\|_{2}$$\leq$ $1.13n^{2}{\bf u}\cdot(\sqrt{1+\epsilon_{b}}+t)\cdot\|L\|_{2}\|U\|_{2}$
$\cdot$ $\frac{1.12}{\sqrt{1-\epsilon_{s}}}\cdot\frac{1}{\sigma_{min}(L)\sigma_{min}(U)}+1.1s_{2}n\sqrt{n}{\bf u}\cdot(\sqrt{1+\epsilon_{b}}+t)\cdot\|L\|_{2}$
$\cdot$ $\frac{1.12}{\sqrt{1-\epsilon_{s}}}\cdot\frac{1}{\sigma_{min}(L)}+t\|L\|_{2}$
$\cdot$ $\frac{1.12}{\sqrt{1-\epsilon_{s}}}\cdot\frac{1}{\sigma_{min}(L)}+1.1n^{2}{\bf u}\cdot\|L\|_{2}\|U\|_{2}\cdot\|\Omega_{1}\|_{2}\|\Omega_{2}\|_{2}$
@@ 4.4.1. Orthogonality
with probability at least $1-p_{f}$. Therefore, with (4.58), we can have
(4.59) $0.8\leq\sigma_{min}(\Omega_{2}\Omega_{1}PXR_{ss}^{-1})\leq\|\Omega_{2}\Omega_{1}PXR_{ss}^{-1}\|_{2}\leq 1.2$
with probability at least $1-p_{f}$. According to Lemma 2.10, when $\Omega_{1}$ and $\Omega_{2}$ embed $PXR_{ss}^{-1}$ and $\Omega_{1}PXR_{ss}^{-1}$, with (2.14) and (2.15), we can have
(4.60) $\frac{4}{5\sqrt{1+\epsilon_{b}}}\leq\sigma_{min}(PXR_{ss}^{-1})\leq\|PXR_{ss}^{-1}\|_{2}\leq\frac{6}{5\sqrt{1-\epsilon_{s}}}$
with probability at least $(1-p_{f})^{2}$. Since $P$ is a permutation matrix, we can easily have
(4.61) $\sigma_{i}(XR_{ss}^{-1})=\sigma_{i}(PXR_{ss}^{-1})$
when $i=1,2,\cdots,n$. Based on (4.59) and (4.61), we can have
(4.62) $\frac{4}{5\sqrt{1+\epsilon_{b}}}\leq\sigma_{min}(XR_{ss}^{-1})\leq\|XR_{ss}^{-1}\|_{2}\leq\frac{6}{5\sqrt{1-\epsilon_{s}}}$
with probability at least $(1-p_{f})^{2}$. Based on (4.5), we can have
(4.63) $Q_{ss}-XR_{ss}^{-1}=E_{QRss}R_{ss}^{-1}.$
With (4.39) and (4.52), when (4.14) is satisfied, we can have
$\|Q_{ss}-XR_{ss}^{-1}\|_{2}$$\leq$ $\|E_{QRss}\|_{2}\|R_{ss}^{-1}\|_{2}$
$\leq$ $1.15n^{2}{\bf u}\cdot(\sqrt{1+\epsilon_{b}}+t)\cdot\|L\|_{2}\|U\|_{2}\|Q_{ss}\|_{2}$
@@
(4.65) $\leq$ $0.06\|Q_{ss}\|_{2}$
with probability at least $1-p_{f}$. Based on Lemma 2.1 and (4.5), we can have
(4.66) $\sigma_{min}(XR_{ss}^{-1})-\|Q_{ss}-XR_{ss}^{-1}\|_{2}\leq\sigma_{min}(Q_{ss})\leq\|Q_{ss}\|_{2}\leq\|XR_{ss}^{-1}\|_{2}+\|Q_{ss}-XR_{ss}^{-1}\|_{2}.$
Based on (4.62) and (4.65), we can get $\|Q_{ss}\|_{2}\leq\frac{6}{5\sqrt{1-\epsilon_{s}}}+0.06\|Q_{ss}\|_{2}$ with probability at least $(1-p_{f})^{2}$. Therefore, it is easy to have
(4.67) $\|Q_{ss}\|_{2}\leq\frac{1.28}{\sqrt{1-\epsilon_{s}}}$
with probability at least $(1-p_{f})^{2}$. With (4.65) and (4.67), we can have $\|Q_{ss}-XR_{ss}^{-1}\|_{2}\leq\frac{0.08}{\sqrt{1-\epsilon_{s}}}$ with probability at least $(1-p_{f})^{2}$. Therefore, based on (4.62), we can have
(4.68) $\sigma_{min}(Q_{ss})\geq\frac{4}{5\sqrt{1+\epsilon_{b}}}-\frac{0.08}{\sqrt{1-\epsilon_{s}}}$
with probability at least $(1-p_{f})^{2}$. With (4.67) and (4.68), we can have
(4.69) $\kappa_{2}(Q_{ss})$$\leq$ $\frac{\|Q_{ss}\|_{2}}{\sigma_{min}(Q_{ss})}$
@@
with probability at least $(1-p_{f})^{2}$. Under (4.10), we can easily get $\frac{1.28}{0.8\cdot\sqrt{\frac{1-\epsilon_{s}}{1+\epsilon_{b}}}-0.08}\leq\frac{1}{8\sqrt{mn{\bf u}+n(n+1){\bf u}}}$. Then based on the theoretical results of [39], we can get
@@
with probability at least $(1-p_{f})^{2}$. Therefore, (4.15) is proved.
4.4.2. Residual
In the following, we focus on the residual. Regarding (4.6) and (4.7), we can have
(4.70) $Q_{1ss}R_{2ss}$ $=$ $(Q_{ss}+E_{3ss})R_{1ss}^{-1}\cdot(R_{1ss}R_{ss}+E_{4ss})$
@@ 4.4.2. Residual
Therefore, with (4.70), we can have
(4.71) $\|Q_{1ss}R_{2ss}-X\|_{F}\leq\|Q_{ss}R_{ss}-X\|_{F}+\|E_{3ss}\|_{F}\|R_{ss}\|_{2}+\|E_{4ss}\|_{F}\|Q_{1ss}\|_{2}.$
With Lemma 2.2, (4.6), (4.7) and (4.71), we can have
(4.72) $\|Q_{1ss}R_{2ss}-X\|_{F}$$\leq$ $\|Q_{ss}R_{ss}-X\|_{F}+\|E_{3ss}\|_{F}\|R_{ss}\|_{2}+\|E_{4ss}\|_{F}\|Q_{1ss}\|_{2}$
$\leq$ $1.1n^{2}{\bf u}\cdot(\|Q_{ss}\|_{2}\|R_{ss}\|_{2}+2\|Q_{1ss}\|_{2}\|R_{1ss}\|_{2}\|R_{ss}\|_{2}).$
For $\|Q_{1ss}R_{2ss}-X\|_{F}$, we can do it in a similar way with that of [34]. Regarding $\|Q_{1ss}\|_{2}$, with theoretical results in [39] and (4.69), we can have
(4.73) $\|Q_{1ss}^{\top}Q_{1ss}-I\|_{F}$$\leq$ $\frac{5}{64}\cdot(8\kappa_{2}(Q_{ss})\sqrt{mn{\bf u}+n(n+1){\bf u}})^{2}$
@@
with probability at least $(1-p_{f})^{2}$. Therefore, we can get
(4.74) $\|Q_{1ss}\|_{2}\leq\sqrt{1+h_{1}}$
with probability at least $(1-p_{f})^{2}$. With corresponding results in [39] and (4.67), we can have
(4.75) $\|R_{1ss}\|_{2}$$\leq$ $\frac{65}{63}\|Q_{ss}\|_{2}$
@@
with probability at least $(1-p_{f})^{2}$. Based on (4.5) and (4.53), we can have
(4.76) $\|Q_{ss}R_{ss}-X\|_{2}$$\geq$ $\|Q_{ss}R_{ss}\|_{2}-\|E_{QRss}\|_{2}$
$\geq$ $\sigma_{min}(Q_{ss})\|R_{ss}\|_{2}-1.1n^{2}{\bf u}\cdot\|Q_{ss}\|_{2}\|R_{ss}\|_{2}.$
@@
(4.77) $\|R_{ss}\|_{2}\leq\frac{\|X\|_{2}}{\sigma_{min}(Q_{ss})-1.1n^{2}{\bf u}\cdot\|Q_{ss}\|_{2}}.$
We put (4.67) and (4.68) into (4.77) and based on (4.12), we can get
(4.78) $\|R_{ss}\|_{2}$$\leq$ $\frac{\|X\|_{2}}{\frac{4}{5\sqrt{1+\epsilon_{b}}}-\frac{0.08}{\sqrt{1-\epsilon_{s}}}-1.1n^{2}{\bf u}\cdot\frac{1.28}{\sqrt{1-\epsilon_{s}}}}$
@@
with probability at least $(1-p_{f})^{2}$. Here, (4.10) guarantees $\frac{4}{5\sqrt{1+\epsilon_{b}}}-\frac{0.11}{\sqrt{1-\epsilon_{s}}}>0$. Therefore, We put (4.67), (4.74), (4.75) and (4.78) into (4.72) and we can have
(4.79) $\|Q_{1ss}R_{2ss}-X\|_{F}$$\leq$ $\|Q_{ss}R_{ss}-X\|_{F}+\|E_{3ss}\|_{F}\|R_{ss}\|_{2}+\|E_{4ss}\|_{F}\|Q_{1ss}\|_{2}$
@@
with probability at least $(1-p_{f})^{2}$.
Regarding $\|Q_{2}^{\top}R_{4}-X\|_{F}$, with (4.8) and (4.9), we can have
(4.80) $Q_{2ss}R_{4ss}$$\leq$ $(Q_{1ss}+E_{7ss})R_{3ss}^{-1}(R_{3ss}R_{2ss}+E_{8ss})$
@@
(4.81) $\|Q_{2ss}R_{4ss}-X\|_{F}\leq\|Q_{1ss}R_{2ss}-X\|_{F}+\|E_{7ss}\|_{F}\|R_{4ss}\|_{2}+\|Q_{2ss}\|_{F}\|E_{8ss}\|_{F}.$
According to the existing results in [39], we can have
(4.82) $\|Q_{2ss}\|_{2}\leq 1.1$
with probability at least $(1-p_{f})^{2}$. Moreover, based on (4.9), we can have
(4.83) $\|R_{4ss}\|_{2}\leq\|R_{3ss}\|_{2}\|R_{2ss}\|_{2}+\|E_{8ss}\|_{2}.$
Similar to (4.83), regarding $\|R_{2ss}\|_{2}$, with (4.7), we can have
(4.84) $\|R_{2ss}\|_{2}\leq\|R_{1ss}\|_{2}\|R_{ss}\|_{2}+\|E_{4ss}\|_{2}.$
With Lemma 2.2, (4.12), (4.75) and (4.78), we can estimate $\|E_{4ss}\|_{F}$ as
(4.85) $\|E_{4ss}\|_{F}$$\leq$ $1.1n^{2}{\bf u}\cdot\|R_{1ss}\|_{2}\|R_{ss}\|_{2}$
@@
with probability at least $(1-p_{f})^{2}$. Therefore, we put (4.75), (4.78) and (4.85) into (4.84) and we can get
(4.86) $\|R_{2ss}\|_{2}$$\leq$ $\|R_{1ss}\|_{2}\|R_{ss}\|_{2}+\|E_{4ss}\|_{2}$
@@
with probability at least $(1-p_{f})^{2}$. Similar to (4.74) and (4.75), we can bound $\|R_{3ss}\|_{2}$ as
(4.87) $\|R_{3ss}\|$$\leq$ $\frac{65}{63}\|Q_{1ss}\|_{2}$
@@
with probability at least $(1-p_{f})^{2}$. Therefore, with Lemma 2.2, (4.12), (4.86) and (4.87), we can estimate $\|E_{8ss}\|_{F}$ as
(4.88) $\|E_{8ss}\|_{F}$$\leq$ $1.1n^{2}{\bf u}\|R_{3ss}\|_{2}\|R_{2ss}\|_{2}$
@@
with probability at least $(1-p_{f})^{2}$. With Lemma 2.2, (4.82) and (4.87), we can bound $\|E_{7ss}\|_{F}$ as
(4.89) $\|E_{7ss}\|_{F}$$\leq$ $1.1n^{2}{\bf u}\|Q_{2ss}\|_{2}\|R_{3ss}\|_{2}$
@@
with probability at least $(1-p_{f})^{2}$. Moreover, with (4.12), we put (4.86)-(4.88) into (4.83) and we can estimate $\|R_{4ss}\|_{2}$ as
(4.90) $\|R_{4ss}\|_{2}$$\leq$ $\|R_{3ss}\|_{2}\|R_{2ss}\|_{2}+\|E_{8ss}\|_{2}$
@@
with probability at least $(1-p_{f})^{2}$. In all, we put (4.79), (4.82), (4.88)-(4.90) into (4.81) and we can get
@@
$\leq$ $[1.79(1+h_{1})+4.63\sqrt{1+h_{1}}+1.41]\cdot\frac{h_{2}}{\sqrt{1-\epsilon_{s}}}\cdot n^{2}{\bf u}\cdot\|X\|_{2}$
with probability at least $(1-p_{f})^{2}$. (4.16) is proved. In all, Theorem 4.1 is proved. ∎
4.5. Proof of Theorem 4.2
Proof.
Regarding the proof of Theorem 4.2, we focus directly on (4.73) and (4.79). When we consider the single-sketching case with $\Omega=\Omega_{2}$, $\Omega_{1}\in\mathbb{R}^{m\times m}$ is the identity matrix $I_{m}$, resulting in $s_{1}=m$ and $\|\Omega_{1}\|_{2}=1$. Consequently, if $\Omega_{1}\in\mathbb{R}^{s_{1}\times m}$ is a $(\epsilon_{1},p_{1},n)$ oblivious $l_{2}$-subspace embedding in $\mathbb{R}^{m}$, we can take $\epsilon_{1}=0$ and $p_{1}=0$. Next, when $\Omega\in\mathbb{R}^{s\times m}$ is a $(\epsilon,p,n)$ oblivious $l_{2}$-subspace embedding in $\mathbb{R}^{s_{1}}$, we let $\epsilon=\epsilon_{b}=\epsilon_{s}$, $p=p_{f}$ and $s=s_{2}$ in (4.73) and (4.79). With these settings, we can derive (4.22) and (4.23). (4.17)-(4.21) are similar to that of (4.10)-(4.14). (4.17) Thus, Theorem 4.2 is proved. ∎
5. Numerical experiments